Determine the phase angle between voltage and current for the following circuit (#2b). Thank you for the help.Determine phase angle between voltage and current.
Answer #1

Here f is required to calculate w.I am doing the procedure by taking w as angular frequency.

Xc=(1/wC) =1/(w*0.0022*10-6) =(4.55*108/w) ohm

R=56000 ohm

Then actual phase angle between generator current and generator voltage is (with voltage lagging current):

θ = tan-1(Xc/R) = tan-1[4.55*108/w)/56000] = tan-1(8125/w)

Put value of w and calculate phase angle by taking tan inverse.